Let's consider any quadrilateral. What is The sum of it's interior angles in degrees?


1
Expert's answer
2021-08-09T11:04:43-0400

A quadrilateral has 4 sides. From this idea, the sum of its interior angles will be equal to the sum of interior angles of n−2 triangles.

where n=4

number of triangles in this case = 4-2= 2 triangles

interior angles of a triangle add up to 180o.

for the two triangles, "2\times 180
^o=360^o"

Thus, the sum of the interior angles of any quadrilateral is 360°.